9a - 2b= C + 4a, for the a
Help please, I realy don't understand.

Answer by tutorcecilia(2152)

9a - 2b= C + 4a
9a-4a-2b=C+4a-4a
5a-2b=C
5a-2b+2b=C+2b
5a=C+2b
5a/5=(C+2b)/5
a=(C+2b)/5
.